# Ceci N’est Pas un Drone: Investigating the Impact of Design Representation on Design Decision Making When Using GenAI

### v1

@@ 1 Introduction
Engineering design is a complex and creative exercise soria_zurita_survey_2017; song_toward_2020; song_when_2021; xu_adaptation_2024; xu_mind_2025 that involves innovative ideation, critical thinking, technical analysis, and iterative problem solving. It also requires designers to make tradeoffs in their search for solutions, balancing between exploration and exploitation in the design space march_exploration_1991; tabeau_exploration_2017; okamoto_proposal_2022. Exploration often requires designers to look for solutions that are original and novel, exemplified by design activities like ideation and brainstorming march_exploration_1991; gupta_interplay_2006. In contrast, exploitation emphasizes the refinement of existing details and making choices towards greater efficiency; this mindset often becomes important during design validation and iterative prototyping march_exploration_1991; gupta_interplay_2006. Artificial Intelligence (AI) technologies present new opportunities for engineers and designers to achieve improved productivity in both exploration and exploitation song_when_2021; xu_adaptation_2024.
@@
Generative AI can substantially support design exploration by rapidly providing an abundance of feasible design solutions, usually significantly more than what would be possible with human effort alone, further highlighting the importance of the subsequent design selection and exploitation process to materialize those innovations oh_deep_2019; zhu_generative_2022; zhu_generative_2023; kim_effect_2023. Engineers and designers, restricted by limited time and resources, will need to use their expertise and judgment to select a smaller group of candidate designs for further validation and prototyping wallace_methods_1995; gembarski_making_2021. Therefore, the ability to judge the quality of designs, especially when presented with numerous options, is critical for the successful implementation of generative design tools hong_generative_2023; chen_how_2025; fang_generative_2025.
Prior studies have shown that the modality in which a design is presented affects people’s perceptions reid_impact_2013; detchprohm_effect_2025; barnawal_evaluation_2017; schulze-meesen_impact_2023; derya_ozcelik_buskermolen_effect_2015, which could potentially affect their judgments. It is common for designers to use visualizations of their designs, such as visual renderings (geometric representations of physical products) or sketches, to aid communications and collaborations (henderson_flexible_1991; haggman_connections_2015; tsai_how_2017; atit_shah_learners_2021). Reid et al. reid_impact_2013 found that different visual representations affect customer subjective preferences and objective measurements of the products, but not their judgments on product function attributes. Also, Detchprohm et al. detchprohm_effect_2025 showed in a study that the quality of visual representations of the designs will not affect the perceived functionality of the product. Studies have also shown that visual representations facilitate better communication and feedback compared to textual information barnawal_evaluation_2017; schulze-meesen_impact_2023. However, results on the differences between visual representations are mixed. One study by Barnawal et al. barnawal_evaluation_2017 suggests that 3D design representations do not have a significant advantage over 2D ones for communicating design concepts, but have higher usability, while another study by Buskermolen et al. derya_ozcelik_buskermolen_effect_2015 found that the motion (stills vs. an animation) and visual quality do not affect concept comprehension, but visual quality can affect the style of communication. Padilla et al. padilla_decision_2018 investigated the effect of different data visualizations on people’s judgments and decision-making when interpreting data, and found that data visualization affects decision-making by invoking different cognitive processes. The complex nature of engineering design tasks requires both objective information delivery and subjective evaluation of the feasibility of design concepts outside of design data. The potential biases introduced by different design representation modalities could weaken designers’ judgment abilities and, therefore, negatively impact the usefulness of generative AI systems in engineering design.
@@
It seems an easy conclusion that humans rely on visualization techniques to do design well, given the ubiquity of design visualization in design practices. Studies have shown that design visualization (e.g., sketches) adds additional information to numerical performance data as a cognitive tool ullman_importance_1990; goldschmidt_dialectics_1991; suwa_what_1997, suggesting that designers have heuristics that enable them to extract non-textual information from design visualization. However, does that necessarily mean designers can make better judgments and design selections with design visualization? Making the design selection process more challenging, AI-generated designs can be "weird" and "unusual", with aesthetics that humans may find unattractive despite performance advantages loos_towards_2022. In the context of engineering design, this raises further questions: Do engineers actually need to see the design visualization to make good judgments and design choices? Or is the numerical performance data, along with proper data visualization, all they need? Do design visualization and geometric rendering carry any additional value to the design decision-making process outside of what design numerical performance data can offer? Research in education has shown that visualization deepens learners’ understanding of the subject and improves learning outcomes as an epistemic object evagorou_role_2015; schoenherr_learning_2024. It is possible that the presence of design visualization and design renderings enables human heuristics and allows engineers and designers to capture features and information that are missed by objective functions or mere numerical performance data. It may be critical to have design renderings and visualization presented for engineers to make accurate judgments. Nevertheless, it is also possible that the presence of design visualization introduces unwanted bias or design fixations, leading to sub-optimal design decisions.
@@
That leads to our central research question: Do different design modalities affect human decision-making behavior and their ability to make optimal selections when presented with AI-generated design solutions? So far, there is insufficient evidence in the design and AI research community to answer this question or to explicitly and systematically examine the impact of design modalities on design decision-making. We hypothesize that different design modalities can affect engineers’ decision-making. We also hypothesize that using a design visualization (geometric representations of physical products) as a representation modality can negatively impact engineers’ ability to identify optimal and novel design solutions.
To test our hypotheses, we designed and conducted a within-subjects experiment across three studies targeting different populations to study whether and how different types of design representations, including design visualizations and numerical design performance data, affect engineers’ design choices and their ability to choose the optimal designs in a list of AI-generated design ideas. These three studies progressively build on one another to address emergent phenomena observed. During a single study, the participants individually considered three UAV (uncrewed aerial vehicles) design problems, where they were asked to select the most optimal drone design from the provided list of design options and briefly explain their choice. For the first study, we targeted drone hobbyists and drone pilots at a U.S. research university. During the first study, we observed that the number of design options provided could affect the participants’ choices. To further investigate the issue, we conducted the second study with STEM students at a U.S. research university through an engineering design related course. For the third study, we recruited aerospace engineers and designers from a large aerospace company in order to examine potential expert-novice differences. Combined, these experiments investigate whether and how different types of design representations affect engineers’ decision-making when working with generative-AI design tools.
@@
Our results confirm both of our hypotheses. We found that different design modalities do affect engineers’ decision-making when using AI-powered generative design tools. Specifically, in this study, designers provided with numerical design performance data only lead to the best ability to select optimal designs, while only seeing design renderings provides marginal help in selecting optimal designs. Unexpectedly, when participants saw both the numerical design performance data and the design rendering, they performed more poorly compared to when they saw the numerical design performance data alone. However, we found that the participants prefer the best-performing designs as long as those designs possess traditional and symmetrical appearances. We also found in this study that the number of design options provided affects people’s ability to choose optimal designs, and a large number of design options can overwhelm engineers and lead to suboptimal design choices.
@@
This paper contributes empirical evidence on how design representation modalities affect designers’ decisions when using AI-powered generative design tools, specifically through:
@@
  • Empirical findings from three studies with drone hobbyists, STEM students, and aerospace professionals, showing that different representation modalities (visual rendering, numerical data, and visual rendering + numerical data) affect design choices with generative design tools.
@@ 2.1 AI in engineering design
In an effort to facilitate solving complex engineering design problems, the engineering design research community has been studying automated design tools since the 1980s maher_hi-rise_1985; smithers_-based_1989. With the rapid development of modern AI and Machine Learning (ML) technologies, a wide variety of studies have investigated the implementation of AI and ML in the engineering design process, including design exploration and concept generation kim_ai_2019; raina_learning_2019; camburn_computer-aided_2020; valdez_framework_2021; zhu_generative_2022; zhu_generative_2023; kim_effect_2023; saadi_generative_2023; khanolkar_mapping_2023; joosten_comparing_2024, design concept evaluation camburn_machine_2020; song_decoding_2022; demirel_human-centered_2024, design optimization sharpe_comparative_2019; nie_topologygan_2021; behzadi_gantl_2021; senhora_machine_2022; wang_generative_2023; maze_diffusion_2023, and prototyping and manufacturing dering_unsupervised_2017; williams_design_2019; qin_research_2022; tercan_machine_2022; kumar_machine_2023. Those implementations of AI in engineering design are believed to improve the efficiency of the design process mirhoseini_graph_2021; yuksel_review_2023, and can even improve the quality of the design solutions joosten_comparing_2024.
@@
More noticeably, a recent focus of the research community is the adoption of AI-powered design tools for design generation oh_deep_2019; chen_inverse_2021; heyrani_nobari_creativegan_2021; regenwetter_deep_2022. AI-powered generative design tools can promptly create a large number of design solutions chen_inverse_2021; heyrani_nobari_creativegan_2021; regenwetter_deep_2022. With the help of AI-powered generative design tools, engineers and designers can search for potential design solutions more efficiently with a larger scope and at a lower cost koch_design_2017; oh_deep_2019; camburn_computer-aided_2020; camburn_machine_2020; dering_unsupervised_2017; byrne_application_2025. In general, generative AI is believed to enhance the design ideation and concept generation process oh_deep_2019; kim_effect_2023. However, the abundance of potential design solutions may pose new challenges for designers when selecting optimal designs.
2.2 Design decision making and judgment ability: key for successful AI implementation
@@
Concept selection and design decision making are an essential part of engineering design wallace_methods_1995; gembarski_making_2021. Restricted by limited time and resources, engineers and designers constantly face trade-offs and design decision-making in the engineering design process division_on_engineering_and_physical_sciences_theoretical_2001; otto_trade-off_1991; nickel_contextual_2024. In order to make informed decisions, designers need to consider different factors and features, and seek a balance between them kalsi_comprehensive_1999; division_on_engineering_and_physical_sciences_theoretical_2001. As a result, the judgment ability and the ability to evaluate designs are critical for the successful implementation of generative design tools hong_generative_2023; chen_how_2025; fang_generative_2025. Experienced engineers and designers can better leverage their domain knowledge to prioritize promising AI suggestions and better utilize this new technology tambe_reskilling_2025. It is important to consider the impact on the judgment abilities of the designers and domain experts when introducing design tools.
2.3 Design representation modalities and the impact on judgment ability
@@
Among the common design modalities, numerical performance data is an intuitive and straightforward approach to conveying design information in an accurate and concise manner, especially when paired with appropriate data visualization abi_akle_graphical_2015; araci_trade-off_2017; cibulski_paved_2020. However, constrained by its textual nature and narrow representation of design features, design information delivered by numerical performance data could be limited in scope (e.g., incomplete or fragmented information on design development and justification cheng_role_2019; mirabito_feature_2024).
@@
In comparison, design visualization (geometric representations of physical products) is another common design modality that is visual-based and more intuitive, and has a more inclusive representation of physical design features (larkin_why_1987; henderson_flexible_1991; ivanov_visualization_2024). Visual design representations contain rich information about the design, especially for the structure of the design, allowing for accurate and comprehensive interpretation (xu_mind_2025; tsai_how_2017). As a result, design visualizations are widely used in different stages of the engineering design process (henderson_flexible_1991; haggman_connections_2015; tsai_how_2017; atit_shah_learners_2021; veisz_computer-aided_2012). It is common practice in the engineering design industry to use visualization of the designs, such as visual renderings or sketches, to aid communications and collaborations (henderson_flexible_1991; haggman_connections_2015; tsai_how_2017; atit_shah_learners_2021). Research has shown the positive effects of using visualization for engineering design practices, including facilitating ideation, communication, and collaboration, and improving shared understanding (suwa_what_1996; mckoy_influence_2001; tversky_what_2002; tversky_sketches_2003; heiser_sketches_2004; macomber_role_2011; worinkeng_can_2013; xu_mind_2025).
@@
However, research has also cautioned the potential risks of overreliance on design visualizations, including design fixation and detriments to design creativity jansson_design_1991; atilola_representing_2015; atilola_effects_2016; amann_fixation_1988; viswanathan_design_2013. These effects can affect designers’ judgment, and hinder their willingness and ability to explore novel solutions suggested by generative AI systems. Furthermore, studies have shown that design modalities could affect people’s perceptions of the design, and, therefore, affect their judgments reid_impact_2013; derya_ozcelik_buskermolen_effect_2015; barnawal_evaluation_2017; schulze-meesen_impact_2023; detchprohm_effect_2025. Judgment biases can also result from the use of particular modalities padilla_decision_2018. All these potential hindrances can weaken designers’ abilities to objectively evaluate designs and select optimal design solutions, and, therefore, negatively impact the development and implementation of generative engineering design tools.

### v2

@@ 1 Introduction
Engineering design is a complex, creative exercise soria_zurita_survey_2017; song_toward_2020; song_when_2021; xu_adaptation_2024; xu_mind_2025 involving innovative ideation, critical thinking, technical analysis, and iterative problem solving. It also requires designers to make tradeoffs in their search for solutions, balancing between exploration and exploitation in the design space march_exploration_1991; tabeau_exploration_2017; okamoto_proposal_2022. Exploration targets original, novel solutions, exemplified by design activities like ideation and brainstorming march_exploration_1991; gupta_interplay_2006, whereas exploitation refines existing concepts and prioritizes efficiency, especially during validation and prototyping march_exploration_1991; gupta_interplay_2006. Artificial Intelligence (AI) technologies present new opportunities for engineers and designers to achieve improved productivity in both exploration and exploitation song_when_2021; xu_adaptation_2024.
@@
Generative AI can substantially support design exploration by rapidly providing an abundance of feasible design solutions, far beyond what would be possible with human effort alone, further highlighting the importance of the subsequent design selection and exploitation process to materialize those innovations oh_deep_2019; zhu_generative_2022; zhu_generative_2023; kim_effect_2023. Engineers and designers, restricted by limited time and resources, must use their expertise and judgment to down-select candidate designs for further validation and prototyping wallace_methods_1995; gembarski_making_2021. Therefore, the ability to judge design quality, especially when presented with numerous options, is critical for the successful implementation of generative design tools hong_generative_2023; chen_how_2025; fang_generative_2025.
In the HCI and information science research community, representation modality has been used to describe the medium and form for information expression, including visual, auditory, haptic, physical, and other media hogan_visual_2017; jansen_opportunities_2015; zong_umwelt_2024; bae_making_2022; demir_rethinking_2025. Drawing from these prior works and considering the engineering context, in the current work, we use representation modality to refer to the form in which design information is presented and delivered. Prior studies have shown that the modality in which a design is presented affects people’s perceptions and subjective preferencesreid_impact_2013; detchprohm_effect_2025; barnawal_evaluation_2017; schulze-meesen_impact_2023; derya_ozcelik_buskermolen_effect_2015; padilla_decision_2018. However, these works mostly focused on graphics information or subjective preferences, rather than on physical engineering designs and objective optimality evaluation. The complex nature of engineering design tasks requires both objective information delivery and subjective evaluation of design concept feasibility outside of design data. Inspired by cognition and neuro-psychology research on judgment and decision-making, we define judgment ability in the current work as a designer’s capacity to make optimal design decisions and selections, after considering available information and relevant circumstances rabin_judgment_2007; capucho_judgment_2011; tversky_judgment_1974. The potential biases introduced by different design representation modalities could weaken designers’ judgment abilities and, therefore, reduce the usefulness of generative AI systems in engineering design.
In this work, we examine generative AI systems from the perspective of what they often afford, particularly in the context of engineering design. Namely, we consider two critical properties of the system in its ability to generate (1) a large set size of the design solutions (GenAI as a design generator), as well as (2) weird or unconventional designs (GenAI for innovative or unhuman-like designs). These attributes characterize the design output from the system, which users directly interact with.
@@
With or without generative AI systems, it seems an easy conclusion that humans rely on visualization techniques to do design well, given the ubiquity of design visualization in design practices. Studies have shown that design visualization (e.g., sketches) adds information to numerical performance data as a cognitive tool ullman_importance_1990; goldschmidt_dialectics_1991; suwa_what_1997; larkin_why_1987; stenning_cognitive_1995, suggesting that designers have heuristics that enable them to extract non-textual information from design visualization. However, AI-generated designs can be "weird" and "unusual", with aesthetics that humans may find unattractive despite performance advantages loos_towards_2022. In the context of engineering design, this raises further questions: Do engineers actually need design visualization to make good judgments and design choices? Or is the numerical performance data, along with proper data visualization, sufficient? Do design visualization and geometric rendering add value to the design decision-making process beyond numerical performance data? Research in education has shown that visualization deepens learners’ understanding of the subject and improves learning outcomes as an epistemic object evagorou_role_2015; schoenherr_learning_2024. It is possible that design visualization and renderings enable human heuristics, allowing engineers and designers to capture features and information that are missed by objective functions or numerical performance data alone. Nevertheless, it is also possible for design visualizations to introduce unwanted bias or design fixation, similar to what prior information visualization and information graphics studies have found, such as exaggerated expectation and illusion of validity dimara_task-based_2020; tversky_judgment_1975; sharp_performance_1988. Such decision biases can ultimately lead to sub-optimal design decisions.
@@
That leads to our central research question: Do different design modalities affect human decision-making behavior and their ability to make optimal selections when presented with AI-generated design solutions? So far, there is insufficient evidence in the design and AI research community to answer this question or to explicitly and systematically examine the impact of design modalities on engineering design decision-making. We hypothesize that (1) different design modalities can affect engineers’ decision-making, and (2) using a design visualization (geometric representations of physical products) as a representation modality can limit engineers’ ability to identify optimal and novel design solutions.
To test our hypotheses, we designed and conducted a within-subjects experiment across two studies targeting different populations, examining whether and how different types of design representations, including design visualizations and numerical design performance data, affect engineers’ design choices and their ability to select optimal designs from a list of AI-generated design ideas.
@@
Our results confirm both of our hypotheses. We found that different design modalities do affect engineers’ decision-making when using AI-powered generative design tools. Specifically, in this study, designers provided with only numerical design performance data had the best ability to select optimal designs, outperforming those who saw both the numerical design performance data and the design rendering. We also found that the participants prefer the best-performing designs as long as those designs possess traditional and symmetrical appearances.
@@
This paper contributes empirical evidence on how design representation modalities affect designers’ decisions when using AI-powered generative engineering design tools, specifically through:
@@
  • Empirical findings from two studies with drone hobbyists and STEM students, showing that different representation modalities (visual rendering, numerical data, and visual rendering + numerical data) affect design choices with generative design tools.
@@
  • Practical implications and recommendations for future development of generative engineering design tools and design comparison tools used with generative AI systems.
As such, the findings of this work could inform the human-AI collaboration loop in engineering design beyond the current GenAI applications, specifically when exploring systems with abundant design options and design novelty.
@@ 2.1 AI in engineering design
To facilitate solving complex engineering design problems, the engineering design research community has been studying automated design tools since the 1980s maher_hi-rise_1985; smithers_-based_1989. With the rapid development of modern AI and Machine Learning (ML) technologies, studies have investigated the implementation of AI and ML in the engineering design process, including design exploration and concept generation kim_ai_2019; raina_learning_2019; camburn_computer-aided_2020; valdez_framework_2021; zhu_generative_2022; zhu_generative_2023; kim_effect_2023; saadi_generative_2023; khanolkar_mapping_2023; joosten_comparing_2024, design concept evaluation camburn_machine_2020; song_decoding_2022; demirel_human-centered_2024, design optimization sharpe_comparative_2019; nie_topologygan_2021; behzadi_gantl_2021; senhora_machine_2022; wang_generative_2023; maze_diffusion_2023, and prototyping and manufacturing dering_unsupervised_2017; williams_design_2019; qin_research_2022; tercan_machine_2022; kumar_machine_2023. Those implementations of AI in engineering design are believed to improve the efficiency of the design process mirhoseini_graph_2021; yuksel_review_2023 and to improve the quality of design solutions joosten_comparing_2024.
@@
More noticeably, recent research has focused on the adoption of AI-powered design tools for design generation oh_deep_2019; chen_inverse_2021; heyrani_nobari_creativegan_2021; regenwetter_deep_2022. AI-powered generative design tools can promptly create large sets of design solutions chen_inverse_2021; heyrani_nobari_creativegan_2021; regenwetter_deep_2022. With the help of AI generative design tools, engineers and designers can search for potential design solutions more efficiently with a larger scope and at a lower cost koch_design_2017; oh_deep_2019; camburn_computer-aided_2020; camburn_machine_2020; dering_unsupervised_2017; byrne_application_2025. In general, generative AI is believed to enhance the design ideation and concept generation process oh_deep_2019; kim_effect_2023. However, the resulting abundance of potential design solutions may pose new challenges for designers when selecting optimal designs.
2.2 Design decision making and judgment ability
@@
Concept selection and design decision making are essential to engineering design wallace_methods_1995; gembarski_making_2021. Restricted by limited time and resources, engineers and designers constantly face trade-offs and design decision-making in the engineering design process division_on_engineering_and_physical_sciences_theoretical_2001; otto_trade-off_1991; nickel_contextual_2024. To make informed decisions, designers must consider various factors and features and seek a balance among them kalsi_comprehensive_1999; division_on_engineering_and_physical_sciences_theoretical_2001. As a result, the judgment ability and the ability to evaluate designs are critical for the successful implementation of generative design tools hong_generative_2023; chen_how_2025; fang_generative_2025. Experienced engineers and designers can leverage their domain knowledge more effectively to prioritize promising AI suggestions and better utilize this new technology tambe_reskilling_2025. It is important to consider the impact on designers’ judgment abilities when introducing design tools.
2.3 Design representation modalities
@@
Among common design modalities, numerical performance data is an intuitive and straightforward approach to conveying design information accurately and concisely, especially when paired with appropriate data visualization abi_akle_graphical_2015; araci_trade-off_2017; cibulski_paved_2020. However, constrained by its textual nature and narrow representation of design features, design information delivered by numerical performance data could be limited in scope (e.g., incomplete or fragmented information on design development and justification cheng_role_2019; mirabito_feature_2024).
@@
In comparison, design visualization (geometric representations of physical products) is another common design modality that is visual-based, more intuitive, and offers a more inclusive representation of physical design features (larkin_why_1987; henderson_flexible_1991; ivanov_visualization_2024). Visual design representations contain rich information about the design, especially its structure, allowing for accurate and comprehensive interpretation (xu_mind_2025; tsai_how_2017; larkin_why_1987). As a result, design visualizations are widely used in different stages of the engineering design process (henderson_flexible_1991; haggman_connections_2015; tsai_how_2017; atit_shah_learners_2021; veisz_computer-aided_2012). It is common practice in the engineering design industry to use visualization of the designs, such as visual renderings or sketches, to aid communications and collaborations (henderson_flexible_1991; haggman_connections_2015; tsai_how_2017; atit_shah_learners_2021). Research has shown the positive effects of using visualization for engineering design practices, including facilitating ideation, communication, and collaboration, and improving shared understanding (suwa_what_1996; mckoy_influence_2001; tversky_what_2002; tversky_sketches_2003; heiser_sketches_2004; macomber_role_2011; worinkeng_can_2013; xu_mind_2025).
@@
However, research has also cautioned about the potential risks of overreliance on design visualizations, including design fixation and detrimental effects on design creativity jansson_design_1991; atilola_representing_2015; atilola_effects_2016; amann_fixation_1988; viswanathan_design_2013. These effects can affect designers’ judgment, and hinder their willingness and ability to explore novel solutions suggested by generative AI systems.
2.4 The impact of design representation modalities on design decision making
Prior studies show that design modalities could influence people’s perceptions of the design reid_impact_2013; derya_ozcelik_buskermolen_effect_2015; barnawal_evaluation_2017; schulze-meesen_impact_2023; detchprohm_effect_2025. Reid et al. reid_impact_2013 found that different visual representations affect customer subjective preferences and objective measurements of the products, but not their judgments on product function attributes. Detchprohm et al. detchprohm_effect_2025 showed that the visual quality will not affect the perceived functionality of the product. These works evaluated representations based on perceived functionality and preference, rather than on objective, quantifiable optimality.
Visual representations can also facilitate communication and feedback compared to textual information barnawal_evaluation_2017; schulze-meesen_impact_2023. However, findings on the differences between visual representations are mixed. Barnawal et al. barnawal_evaluation_2017 suggested that 3D design representations did not outperform 2D ones for communicating design concepts, but they improved usability, while Buskermolen et al. derya_ozcelik_buskermolen_effect_2015 found that the motion (stills vs. animation) and visual quality did not affect concept comprehension, but visual quality can affect the nature and quality of user feedback.
Moreover, Padilla et al. padilla_decision_2018 showed that data graphics influence judgments by invoking different cognitive processes. However, their work focused on abstract data graphics and data visualizations, rather than the visualizations of physical engineering designs. Chen et al. chen_how_2025 explored textual (ChatGPT) and image-based (Midjourney) generative systems in conceptual design, and found that GenAI can support designers in the problem definition and idea generation stages, but not the idea selection and evaluation stage. The GenAI systems examined in their work emphasize early-stage ideation, whereas generative tools built for physical engineering design must also support design selection under physical constraints, engineering feasibility, and objective optimality.
Collectively, these studies suggest different design modalities can steer how designers perceive and evaluate designs. However, they provide limited evidence on how design modalities affect the designers’ objective evaluation of the optimality of physical engineering designs, especially when GenAI produces large sets of unconventional solutions. Our work addresses this gap by providing empirical evidence on the impact of design modality on designers’ ability to identify objectively best-performing designs generated by an AI system.